Last year, CKS had sales of $8B, EBIT of $50m, Net Working Capital of $108m, and capital expenditures matching its depreciation. The year before that, BKS’s Net Working Capital was $120m. CKS has 60m shares outstanding, its effective tax rate is 20% and its debt ratio is 30%. Investors require a 5% (unadjusted) yield on the debt and 9% return on equity.

The discount rate for the FCFF model is_________ %. (Provide your answer in percent, rounded to two decimals, omitting the % sign.)

The share price estimate, if CKS will continue to operate infinitely with a constant free cash flow to the firm growth rate of 6%, is _______
Answer #1

Discount rate, r = WACC = wd x rd x (1 - tax) + we x re = 30% x 5% x (1 - 20%) + 70% x 9% = 7.50%

FCFF = EBIT x (1 - tax) + Depreciation - Change in NWC - Capex

= 50 x (1 - 20%) - (108 - 120)

= 52 million

Value of firm = FCFF x (1 + g) / (r - g) = 52 x (1 + 6%) / (7.50% - 6%) = $3,674.67 million

Equity = 70% x 3,674.67 = $2,572.27 million

Stock Price = 2,572.27 / 60 = $42.87